solve the following initial value problem

II. Resuelve el problema de valore inicial. -ydx + (x + xy)dy = 0 y(1) = 1
Answer #1

We have, -ydx + (x + xy) dy = 0 or rydrtudy say.dy =0 d x.dy- y.da tray dy = 0 kdy - yon = sary dy OY, rept - Lex say dy 24 2The answer sheet has three pages.it is the first page= - lnly) + 9 2 where an integrating constant Janda = pent and nti dr In(all te 거 -1/2 je - enly I to -1/2 -1/2 OY. = -lolltesecond pagec = -2 we get Petting the value of c=-2 in ① -2.5 7 = -lo/11 - 2 Or, 2. √ √ = enly) +2 Therefore, the solection of the givenThird/last page